Qd = 50- 8P
Qs = -17.5 + 10P
a) At the equilibrium price calculate the consumer’ and producers’ surplus.
b) If a price floor of shs 5 is imposed on good X in this market
i) what is the new consumers’ surplus and by how much has it changed?
ii) what is the new producers’ surplus and by how much has it changed?
c) Calculate the transferred surplus. And state from who to whom is it.
d) Calculate the deadweight loss of the price floor
1
Expert's answer
2018-11-09T15:40:09-0500

Qd = 50- 8P

Qs = -17.5 + 10P

a) In equilibrium Qd = Qs, so:

50 - 8P = -17.5 + 10P,

18P = 67.5,

P = 3.75,

Q = 50 - 8*3.75 = 20 units.

The consumer surplus is:

CS = 0.5*(6.25 - 3.75)*20 = 25,

The producer surplus is:

PS = 0.5*(3.75 - 1.75)*20 = 20.

b) If a price floor of shs 5 is imposed on good X in this market, then P = 5, Q = Qd = 50 - 8*5 = 10.

i) the new consumer surplus is:

CS = 0.5*(6.25 - 5)*10 = 6.25, so it has decreased by 18.75.

ii) the new producer surplus is:

PS = 0.5*(2.75 - 1.75)*10 + (5 - 2.75)*10 = 5 + 22.5 = 27.5, so it has increased by 7.5.

c) The transferred surplus is 18.75 - 6.25 or 10*(5 - 3.75) = 12.5. It is transferred from consumers to producers.

d) The deadweight loss of the price floor is the difference between total surpluses before and after imposing the price floor, so:

DWL = (25 + 20) - (6.25 + 27.5) = 45 - 33.75 = 11.25.